OpenStudy (anonymous):

help with negative exponent fractions?! (x+3)^(-1/3)-2(x+3)^(-1/3)=3

OpenStudy (anonymous):

would it be u=(x+3^/1/3)^2?

hero (hero):

Negative exponent means "inverse" not "negative"

hero (hero):

For example \[x^{-a} = \frac{1}{x^a}\]

hero (hero):

In this case \[(x+3)^{-1/3} = \frac{1}{(x+3)^{1/3}}\]

hero (hero):

Which becomes \[\frac{1}{\sqrt[3]{x+3}}\]

hero (hero):

So if you have x to the `negative` a, the negative exponent indicates the `inverse` of x to the a

hero (hero):

And just to remember this, you can basically read \(x^{-a}\) as `the inverse of x to the a` rather than say `x to the negative a`